OpenStudy (anonymous):

Hypothesis testing - A car insurance company claims that it will save new customers 15% of what they pay for their current plans with other insurance companies. In a random sample of 30 new customers, the average amount saved was 8% with a standard deviation of 2%.

OpenStudy (perl):

this needs more information. whats the question

jimthompson5910 (jim_thompson5910):

I'm assuming you're using the default 5% significance level Find the z-score to get z = (xbar - mu)/(sigma/sqrt(n)) z = (0.08 - 0.15)/(0.02/sqrt(30)) z = -19.1702895 Since this standard normal z score is definitely outside the interval from -1.96 to 1.96, this means that this score lies in the rejection region. So you would reject the null hypothesis (that mu = 0.15) and you would be forced to accept the alternative hypothesis (that mu doesn't equal 0.15)

OpenStudy (anonymous):

@jim_thompson5910 can you explain the steps to me please ?

OpenStudy (perl):

what is your question

jimthompson5910 (jim_thompson5910):

we're told that "A car insurance company claims that it will save new customers 15%", so the company claims that the average savings is 15% or 0.15 therefore, the population mean mu is mu = 0.15 ------------------------------- because we have a random sample of 30 people, we know that n = 30 -------------------------------- it then says that for that sample, the "average amount saved was 8%", so the sample mean xbar is xbar = 0.08 -------------------------------- the standard deviation is sigma = 0.02 since it says " with a standard deviation of 2%"

jimthompson5910 (jim_thompson5910):

So that sets up all the given variables and their corresponding values

jimthompson5910 (jim_thompson5910):

you then take the raw score of 0.08 and convert it to a z-score like so z = (xbar - mu)/(sigma/sqrt(n)) z = (0.08 - 0.15)/(0.02/sqrt(30)) z = -19.1702895

OpenStudy (anonymous):

so i take the absolute value of that and since | -19 | = 19 19 > 1.96 i reject the null hypothesis with 95% certainty right ?

jimthompson5910 (jim_thompson5910):

The two hypotheses are Null Hypothesis: Ho: mu = 0.15 Alternative Hypothesis Ha: mu =/= 0.15 So that means we're doing a two tailed test. At the default significance level of 5%, the area in the tails is 0.05/2 = 0.025 You would then use a calculator or a table to find that P(Z < -1.96) = 0.025 or P(Z > 1.96) = 0.025 This means that the acceptance region is from z = -1.96 to z = 1.96 If you get a z-score in this region, then you fail to reject the null hypothesis (and must accept it). If not, then you reject the null hypothesis

jimthompson5910 (jim_thompson5910):

So because we got a z-score that's outside the interval from z = -1.96 to z = 1.96, we reject the null hypothesis

OpenStudy (anonymous):

thank you so much ! i appreciate it

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (perl):

and the population mean x bar is equal to population mean